OpenStudy (anonymous):

Using the graph below, calculate the average rate of change for f(x) from x = 0 to x = 2.

OpenStudy (anonymous):

exponential function going through points 0, −2 and 2, 6 x = −4 x = −2 x = 2 x = 4

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

m = y2-y1/x2-x1 m = 6+2/2-0 m = 8/2 m = 4